Calculate the pH of each of the solutions and the change in pH to 0.01 pH units caused by adding 10.0 mL of 3.59-M HCl to 450. mL of each of the following solutions. Change is defined as final minus initial, so if the pH drops upon mixing the change is negative.

a) water

pH before mixing = 7.00

pH after mixing= 1.10

pH change = -5.90

b) 0.103 M C2H3O2^1-

pH before mixing =

pH after mixing=

pH change =

c) 0.103 M HC2H3O2

pH before mixing = 2.87

pH after mixing=

pH change =

d) a buffer solution that is 0.103 M in each C2H3O2^1- and HC2H3O2

pH before mixing =

pH after mixing=

pH change =
Answer #1

pH of each of the solutions and the change in pH caused by adding 10.0 mL of 3.59 M HCl to 450 mL of each of thesolutions.

b) 0.103 M CH3COO-

pH before mixing :Acetate is a weak base.

A- is a weak base, Kb for it = Kw / Ka (CH3COOH) = 10-14 / (1.8*10-5) = 5.56*10-10

CH3COO-(aq) + H2O (l) \rightleftharpoons  CH3COOH (aq) + OH- (aq)

Kb = (CH3COOH ) (OH- ) / (CH3COO-)

[OH–] = (Kb*   ( CH3COO-) ) ^1/2 = (5.56*10-10*  0.103) ^1/2 = 7.57 *10-6 M

pH = 14 -pOH = 14 - (-log[7.57*10-6 ] = 8.88

pH after mixing :

CH3COO-(aq) + H3O+ (aq) \rightleftharpoons  CH3COOH (aq) + H2O (l)

Any such solution containing comparable amounts of a weak acid , and its conjugate
weak base, is act as buffer. So, pH can be calculated following equation :

pH = pKa + log ( ( A-) / (HA ) )

( CH3COO-) = (0.103 M*450 ml - 3.59 M*10 ml )/ 460 ml = 0.023 M

(CH3COOH ) =  3.59 m*10 ml / 460 ml = 0.078 M

pH = 4.74 + log ( (0.023) / 0.078 ) ) = 4.21

pH change = -4.67

c) 0.103 M CH3COOH

pH before mixing = 2.87

pH after mixing : adding 10.0 mL of 3.59 M HCl will increase H3O+ concentration :

pH before mixing : 2.87 ; so,  H3O+ = 10-2.87 M

( H3O+ ) after mixing =  10-2.87 M* 450 ml + 10 ml * 3.59 M / 460 ml = 0.079

pH = -log (H3O+ ) = 1.10

pH change = - 1.77

d) a buffer solution that is 0.103 M in each CH3COO- and CH3COOH

pH before mixing =

pH can be calculated following equation :

pH = pKa + log ( ( A-) / (HA ) )

pH = 4.74 + log ( (0.103) / 0.103 ) ) = 4.74

pH after mixing : we have following reaction ,

CH3COO-(aq) + H3O+ (aq) \rightleftharpoons  CH3COOH (aq) + H2O (l)

New, concentrations : ( CH3COO-) = (0.103 M*450 ml - 3.59 M*10 ml )/ 460 ml = 0.023 M

(CH3COOH ) = ( 0.103 M*450 ml+ 3.59 m*10 ml) / 460 ml = 0.179 M

pH = 4.74 + log ( (0.0.023) / 0.179 ) ) = 3.85

pH change = - 0.89